# GeoRanker: Distance-Aware Ranking for Worldwide Image Geolocalization

### v1

@@ Abstract
Worldwide image geolocalization—the task of predicting GPS coordinates from images taken anywhere on Earth—poses a fundamental challenge due to the vast diversity in visual content across regions. While recent approaches adopt a two-stage pipeline of retrieving candidates and selecting the best match, they typically rely on simplistic similarity heuristics and point-wise supervision, failing to model spatial relationships among candidates. In this paper, we propose GeoRanker, a distance-aware ranking framework that leverages large vision-language models to jointly encode query–candidate interactions and predict geographic proximity. In addition, we introduce a multi-order distance loss that ranks both absolute and relative distances, enabling the model to reason over structured spatial relationships. To support this, we curate GeoRanking, the first dataset explicitly designed for geographic ranking tasks with multimodal candidate information. GeoRanker achieves state-of-the-art results on two well-established benchmarks (IM2GPS3K and YFCC4K), significantly outperforming current best methods.
@@ 1 Introduction
Figure 1: Accuracy at 1km error threshold for G3 (Current SOTA) vs. the best candidate within top-k retrieved results.
@@
Recent state-of-the-art approaches vivanco2023geoclip ; haas2024pigeon ; zhou2024img2loc ; jia2024g3 typically follow a two-stage pipeline: (1) retrieving and generating a set of candidates based on a global database, and (2) selecting the top match as the predicted geolocation. As shown in Figure 1, although the current SOTA model (G3 jia2024g3 , which ranks candidates via GPS location-image embedding similarity) achieves 16.7% accuracy at 1km error threshold on IM2GPS3K, better candidates often exist among the top-k retrieved results. This suggests that one can retrieve reasonably high-quality candidates, yet the final prediction accuracy hinges on the second stage—the model’s ability to compare spatial relevance and select the most plausible candidate. Currently, the candidate selection is often limited by naïve heuristics such as cosine similarity vivanco2023geoclip ; jia2024g3 , which generally encode the query image and candidates independently, without modeling their mutual spatial relationships or allowing rich interactions. As a result, these methods frequently struggle to distinguish between visually similar yet geographically distant scenes. Furthermore, existing training objectives primarily focus on point-wise similarity between individual images and locations vivanco2023geoclip ; haas2024pigeon ; seo2018cplanet , overlooking the rich spatial relationships among candidates—such as the spatial dependence (i.e., Tobler’s First Law of Geography miller2004tobler ) and relative distances between them—which are crucial for geolocalization.
@@ 2 Related Work
Image Geolocalization. Worldwide geolocalization lies at the intersection of geography and computer vision, and is a core topic in GeoAI mai2024opportunities ; janowicz2020geoai ; li2024georeasoner and spatial data mining wang2020deep ; liang2025foundation . Existing methods for worldwide geolocalization can be grouped into three main categories: classification-based, retrieval-based, and RAG-based approaches. (1) Classification-based methods weyand2016planet ; seo2018cplanet ; pramanick2022world ; clark2023we ; haas2024pigeon approach the task by partitioning the Earth’s surface into discrete Geo-grids and predicting the index of the grid that contains the image location. The final output is typically the center coordinate of the predicted grid. While these methods offer scalability, they may incur large errors when the true location lies far from the grid center, even if the grid prediction is correct. (2) Retrieval-based methods zhu2022transgeo ; lin2022joint ; zhang2023cross ; workman2015wide ; liu2019lending ; zhu2021vigor cast geolocalization as a similarity search problem. These methods either use a database of geotagged images yang2021cross ; zhu2022transgeo ; tian2017cross ; shi2020looking ; zhu2023r2former or a gallery of GPS points vivanco2023geoclip , returning the coordinates of the most similar entries to the input image as the prediction. However, these methods fail to capture the complex spatial relationships between the query image and candidate locations, making it difficult to reliably identify the most accurate match from the candidate pool. (3) RAG-based methods zhou2024img2loc ; jia2024g3 first retrieve a set of candidate locations similar to the query image from the database, then construct a prompt that integrates both the query and candidate information. This prompt is passed to an LVLM to generate a plausible GPS location. In contrast to the above approaches, our proposed Distance-Aware Ranking framework, GeoRanker, focuses specifically on the candidate ranking stage. By explicitly modeling the complex spatial relationships between the query image and candidate geographic entities, an aspect overlooked by prior work, our approach offers more reliable candidate selection and leads to improved geolocalization performance at the global scale.
@@
Learning to Rank. Learning-to-rank cao2007learning (LTR) is a fundamental research direction in information retrieval liu2009learning and recommender systems karatzoglou2013learning , primarily used to train ranking models that refine the order of retrieved candidates based on a given query kabir2024survey . Depending on their modeling and optimization strategies, LTR methods are typically categorized into three types: pointwise, pairwise, and listwise approaches. Pointwise methods bell2018title ; dadaneh2020arsm take the ranking task as a regression or classification problem by assigning a relevance score or label to each query–candidate pair independently. This approach is simple and straightforward, yet overlooks the relative relationships among candidates. Pairwise methods tagami2013ctr ; cerrato2020fair ; jia2021pairrank model the relative preferences between pairs of candidates for the same query. Pairwise methods encourage the model to assign higher scores to positive candidates while penalizing negative ones, learning the relative preferences between different candidates. Listwise methods burges2010ranknet ; xia2008listwise ; cao2007learning ; xu2007adarank can be seen as an extension of pairwise approaches, as they consider the entire list of candidates associated with a query and optimize a loss function that directly reflects the overall quality of the ranking. Our approach builds on the LTR foundation but adapts it to spatial ranking by explicitly modeling and optimizing distance-aware relationships between candidates.
@@ 3.4 Inference
During inference, GeoRanker integrates both retrieved candidates from a database and generated candidates from a LVLM, following prior work zhou2024img2loc ; jia2024g3 . Given a query image $q$, we first retrieve a set of candidates $\mathcal{C}_{\text{r}}$ and collect contextual negative samples $\mathcal{C}_{\text{neg}}$. Simultaneously, the query $q$ is passed through an LVLM to generate a new set of candidates $\mathcal{C}_{\text{g}}$, referred to as generated candidates. The prompt for generating candidates is detailed in Appendix A. We then form query–candidate pairs by combining $q$ with each candidate $c\in\mathcal{C}_{\text{r}}\cup\mathcal{C}_{\text{g}}$, and feed these inputs into GeoRanker to obtain a set of distance scores:
$s_{c}=\text{GeoRanker}(q,c),\quad\forall c\in\mathcal{C}_{\text{r}}\cup\mathcal{C}_{\text{g}}$ (6)
Finally, we select the candidate with the highest score and use its GPS coordinates as the prediction:
@@ 4.2 Main Results
As shown in Table 1, GeoRanker achieves state-of-the-art performance across all evaluation thresholds. For example, on IM2GPS3K, it improves the most challenging street-level accuracy by 12.9% over the best baseline G3, and on YFCC4K, it achieves an 37.3% relative gain at the same threshold. Among the baselines, GeoCLIP vivanco2023geoclip , Img2Loc zhou2024img2loc , PIGEON haas2024pigeon , and G3 jia2024g3 exhibit relatively strong performance due to classification-based methods are limited by systemic biases from fixed candidate grids. Compared to these stronger baselines, our method GeoRanker achieves superior results by explicitly modeling the spatial relationship between each query–candidate pair using a multi-order distance optimization objective. This enables the model to accurately identify the geographically closest candidate as the prediction, further enhancing geolocalization accuracy. In summary, our approach achieves state-of-the-art performance across all datasets and metrics, demonstrating its effectiveness and superiority. Furthermore, Appendix E presents representative examples across various error thresholds to offer intuitive insights into the distribution of query images at different localization accuracies.
@@ Methods
[…]
@@
Table 2 presents an ablation study on the IM2GPS3K dataset, and the results for YFCC4K are illustrated in Appendix F. From Table 2 we draw several key insights: (1) All components in our framework contribute positively to the final performance, demonstrating the effectiveness of our design. (2) Comparing our full model with the variant without second-order distance loss ($\mathcal{L}_{\text{PL}}^{(2)}$), we observe more substantial improvements at coarse-grained levels (e.g., country and continent). This highlights the benefit of modeling second-order spatial relationships among candidates, which enables finer-grained ranking and enhances geolocalization accuracy. (3) Removing any of the modality-aware prompt components—such as negative candidates ($\mathcal{C}_{\text{neg}}$), textual descriptions ($c_{m}^{\text{text}}$), or image data ($c_{m}^{\text{img}}$)—leads to performance drops, confirming that incorporating multi-modal cues into the prompt is beneficial. Among these, visual information yields the most significant gain, underscoring the importance of image semantics. (4) Finally, the variant without generated candidates ($\mathcal{C}_{g}$) underperforms our method, showing that generated candidates provide complementary value. This is especially important in scenarios where the retrieval database lacks relevant examples, and generation can introduce novel, informative candidates that enhance the overall candidate pool.
@@ 4.4 Hyperparameter Analysis
Figure 3: Hyperparameter analysis at the region level on IM2GPS3K. Trends observed at the region level are representative across different geographic levels. Results for all hyperparameters across all levels can be found in Appendix G.
To better understand the impact of key hyperparameters, we conduct a systematic ablation study by varying each at a time while keeping others fixed. The hyperparameters considered include: (1) Number of retrieval candidates in training $\rvert\mathcal{C}_{{rc}}\rvert$. (2) Number of retrieval candidates during inference $\rvert\mathcal{C}_{{r}}\rvert$. (3) Number of generated candidates during inference $\rvert\mathcal{C}_{{g}}\rvert$. (4) Weighting coefficient $\lambda$. (5) Number of top elements involved in the first-order loss ($K^{(1)})$. Unless otherwise specified, we use the following default values: $\rvert\mathcal{C}_{{rc}}\rvert=5$, $\rvert\mathcal{C}_{{r}}\rvert=10$, $\rvert\mathcal{C}_{{g}}\rvert=5$, $\lambda_{1}=0.7$, and $K^{(1)}=1$. Empirically, we find that the trends of each hyperparameter remain largely consistent across all levels, indicating the stability and robustness of our model under varying geolocalization granularities. For clarity and brevity, we present results at the region level as a representative example in Figure 3, while full results for all levels are provided in Appendix G.
@@ 4.7 Impact of Backbone Scale
To investigate the impact of backbone model scale on performance, we conduct experiments using llava-onevision li2024llava (0.5B) and Qwen2-VL models wang2024qwen2 with 2B and 7B parameters. As shown in Figure 6, and results across all geographic levels in Appendix H, GeoRanker’s performance consistently improves as the backbone model size increases on both IM2GPS3K and YFCC4K. These results indicate that GeoRanker benefits from more powerful LVLM backbones and follows the scaling law, suggesting that its upper-bound performance can be further improved with larger models.
@@ Acknowledgements
The authors would like to thank Leitian Tao for the valuable feedback on the work.
@@ Appendix D Baseline Method Details
[…]
@@
  • CPlaNet [15]. CPlaNet follows PlaNet and propose combinatorial partitioning, which generates fine-grained output class by intersecting larger partitions.
@@
  • Img2Loc [10]. Img2Loc advances geolocalization by integrating a RAG pipeline. It first retrieves visually similar candidates, then formulates a prompt incorporating these candidates’ coordinates, guiding a vision-language model to generate a final prediction.
@@
  • G3 [14]. G3 proposes a three-stage framework comprising Geo-alignment, Geo-diversification, and Geo-verification. Geo-alignment aligns GPS coordinates, textual descriptions, and visual data into a unified multi-modal representation to strengthen retrieval capabilities. Subsequently, Geo-diversification and Geo-verification are integrated within a RAG framework to robustly generate and select candidate geolocations.
@@ Appendix E Query Images with Different Error Thresholds
Appendix F Complete experimental results on ablation study

### v3

# GeoRanker: Distance‑Aware Ranking for Worldwide Image Geolocalization

@@ Abstract
Worldwide image geolocalization—the task of predicting GPS coordinates from images taken anywhere on Earth—poses a fundamental challenge due to the vast diversity in visual content across regions. While recent approaches adopt a two-stage pipeline of retrieving candidates and selecting the best match, they typically rely on simplistic similarity heuristics and point-wise supervision, failing to model spatial relationships among candidates. In this paper, we propose GeoRanker, a distance-aware ranking framework that leverages large vision-language models to jointly encode query–candidate interactions and predict geographic proximity. In addition, we introduce a multi-order distance loss that ranks both absolute and relative distances, enabling the model to reason over structured spatial relationships. To support this, we curate GeoRanking, the first dataset explicitly designed for geographic ranking tasks with multimodal candidate information. GeoRanker achieves state-of-the-art results on two well-established benchmarks (IM2GPS3K and YFCC4K), significantly outperforming current best methods. We also release our code, checkpoint, and dataset online [^1] https://github.com/Applied-Machine-Learning-Lab/GeoRanker for ease of reproduction.
@@ 1 Introduction
Figure 1: Accuracy at 1km error threshold for G3 (Current SOTA) vs. the best candidate within top‑k retrieved results.
@@
Recent state-of-the-art approaches vivanco2023geoclip ; haas2024pigeon ; zhou2024img2loc ; jia2024g3 typically follow a two-stage pipeline: (1) retrieving and generating a set of candidates based on a global database, and (2) selecting the top match as the predicted geolocation. As shown in Figure 1, although the current SOTA model (G3 jia2024g3 , which ranks candidates via GPS location-image embedding similarity) achieves 16.7% accuracy at the 1km error threshold on IM2GPS3K, better candidates often exist among the top‑k retrieved results. This suggests that one can retrieve reasonably high-quality candidates, yet the final prediction accuracy hinges on the second stage—the model’s ability to compare spatial relevance and select the most plausible candidate. Currently, the candidate selection is often limited by naïve heuristics such as cosine similarity vivanco2023geoclip ; jia2024g3 , which generally encode the query image and candidates independently, without modeling their mutual spatial relationships or allowing rich interactions. As a result, these methods frequently struggle to distinguish between visually similar yet geographically distant scenes. Furthermore, existing training objectives primarily focus on point-wise similarity between individual images and locations vivanco2023geoclip ; haas2024pigeon ; seo2018cplanet , overlooking the rich spatial relationships among candidates—such as the spatial dependence (i.e., Tobler’s First Law of Geography miller2004tobler ) and relative distances between them—which are crucial for geolocalization.
@@ 2 Related Work
Image Geolocalization. Worldwide geolocalization sarkar2024gomaa ; astruc2024openstreetview ; dufour2025around ; dou2024gaga lies at the intersection of geography and computer vision, and is a core topic in GeoAI mai2024opportunities ; janowicz2020geoai ; li2024georeasoner and spatial data mining wang2020deep ; liang2025foundation . Existing methods for worldwide geolocalization can be grouped into three main categories: classification-based, retrieval-based, and RAG-based approaches. (1) Classification-based methods weyand2016planet ; seo2018cplanet ; pramanick2022world ; clark2023we ; haas2024pigeon approach the task by partitioning the Earth’s surface into discrete Geo-grids and predicting the index of the grid that contains the image location. The final output is typically the center coordinate of the predicted grid. While these methods offer scalability, they may incur large errors when the true location lies far from the grid center, even if the grid prediction is correct. (2) Retrieval-based methods zhu2022transgeo ; lin2022joint ; zhang2023cross ; workman2015wide ; liu2019lending ; zhu2021vigor cast geolocalization as a similarity search problem. These methods either use a database of geotagged images yang2021cross ; zhu2022transgeo ; tian2017cross ; shi2020looking ; zhu2023r2former or a gallery of GPS points vivanco2023geoclip , returning the coordinates of the most similar entries to the input image as the prediction. However, these methods fail to capture the complex spatial relationships between the query image and candidate locations, making it difficult to reliably identify the most accurate match from the candidate pool. (3) RAG-based methods zhou2024img2loc ; jia2024g3 first retrieve a set of candidate locations similar to the query image from the database, then construct a prompt that integrates both the query and candidate information. This prompt is passed to an LVLM to generate a plausible GPS location. In contrast to the above approaches, our proposed Distance-Aware Ranking framework, GeoRanker, focuses specifically on the candidate ranking stage. By explicitly modeling the complex spatial relationships between the query image and candidate geographic entities, an aspect overlooked by prior work, our approach offers more reliable candidate selection and leads to improved geolocalization performance at the global scale.
@@
Learning to Rank. Learning-to-rank cao2007learning (LTR) is a fundamental research direction in information retrieval liu2009learning and recommender systems karatzoglou2013learning , primarily used to train ranking models that refine the order of retrieved candidates based on a given query kabir2024survey . Depending on their modeling and optimization strategies, LTR methods are typically categorized into three types: pointwise, pairwise, and listwise approaches. Pointwise methods bell2018title ; dadaneh2020arsm take the ranking task as a regression or classification problem by assigning a relevance score or label to each query–candidate pair independently. This approach is simple and straightforward, yet it overlooks the relative relationships among candidates. Pairwise methods tagami2013ctr ; cerrato2020fair ; jia2021pairrank model the relative preferences between pairs of candidates for the same query. Pairwise methods encourage the model to assign higher scores to positive candidates while penalizing negative ones, learning the relative preferences between different candidates. Listwise methods burges2010ranknet ; xia2008listwise ; cao2007learning ; xu2007adarank can be seen as an extension of pairwise approaches, as they consider the entire list of candidates associated with a query and optimize a loss function that directly reflects the overall quality of the ranking. Our approach builds on the LTR foundation but adapts it to spatial ranking by explicitly modeling and optimizing distance-aware relationships between candidates.
@@ 3.4 Inference
During inference, GeoRanker integrates both retrieved candidates from a database and generated candidates from an LVLM, following prior work zhou2024img2loc ; jia2024g3 . Given a query image $q$, we first retrieve a set of candidates $\mathcal{C}_{\text{r}}$ and collect contextual negative samples $\mathcal{C}_{\text{neg}}$. Simultaneously, the query $q$ is passed through an LVLM to generate a new set of candidates $\mathcal{C}_{\text{g}}$, referred to as generated candidates. The prompt for generating candidates is detailed in Appendix A. We then form query–candidate pairs by combining $q$ with each candidate $c\in\mathcal{C}_{\text{r}}\cup\mathcal{C}_{\text{g}}$, and feed these inputs into GeoRanker to obtain a set of distance scores: $s_{c}=\text{GeoRanker}(q,c,\mathcal{C}_{\text{neg}}),\quad\forall c\in\mathcal{C}_{\text{r}}\cup\mathcal{C}_{\text{g}}$. Finally, we select the candidate with the highest score and use its GPS coordinates as the prediction:
@@ 4.2 Main Results
As shown in Table 1, GeoRanker achieves state-of-the-art performance across all evaluation thresholds. For example, on IM2GPS3K, it improves the most challenging street-level accuracy by 12.9% over the best baseline G3, and on YFCC4K, it achieves an 37.3% relative gain at the same threshold. Among the baselines, GeoCLIP vivanco2023geoclip , Img2Loc zhou2024img2loc , PIGEON haas2024pigeon , and G3 jia2024g3 exhibit relatively strong performance due to classification-based methods are limited by systemic biases from fixed candidate grids. Compared to these stronger baselines, our method GeoRanker achieves superior results by explicitly modeling the spatial relationship between each query–candidate pair using a multi-order distance optimization objective. This enables the model to accurately identify the geographically closest candidate as the prediction, further enhancing geolocalization accuracy. In summary, our approach achieves state-of-the-art performance across all datasets and metrics, demonstrating its effectiveness and superiority. Furthermore, Appendix E presents representative examples across various error thresholds to offer intuitive insights into the distribution of query images at different localization accuracies. Appendix F validates the ranking capability of GeoRanker using standard learning-to-rank metrics.
@@ Methods
$\text{w/o}\mathcal{L}_{\text{PL}}^{(2)}$ 18.48 44.61 60.96 75.61 88.28
$\text{w/o} \mathcal{C}_{\text{neg}}$ 17.35 44.51 60.82 76.37 88.28
$\text{w/o} c_{m}^{\text{text}}$ 18.02 43.91 60.19 76.61 88.62
$\text{w/o} c_{m}^{\text{img}}$ 15.58 41.77 59.15 75.40 88.35
$\text{w/o} \mathcal{C}_{g}$ 18.21 43.47 59.69 75.47 88.75
@@
Table 2 presents an ablation study on the IM2GPS3K dataset, and the results for YFCC4K are illustrated in Appendix G. From Table 2 we draw several key insights: (1) All components in our framework contribute positively to the final performance, demonstrating the effectiveness of our design. (2) Comparing our full model with the variant without second-order distance loss ($\mathcal{L}_{\text{PL}}^{(2)}$), we observe more substantial improvements at coarse-grained levels (e.g., country and continent). This highlights the benefit of modeling second-order spatial relationships among candidates, which enables finer-grained ranking and enhances geolocalization accuracy. (3) Removing any of the modality-aware prompt components—such as negative candidates ($\mathcal{C}_{\text{neg}}$), textual descriptions ($c_{m}^{\text{text}}$), or image data ($c_{m}^{\text{img}}$)—leads to performance drops, confirming that incorporating multi-modal cues into the prompt is beneficial. Among these, visual information yields the most significant gain, underscoring the importance of image semantics. (4) Finally, the variant without generated candidates ($\mathcal{C}_{g}$) underperforms our method, showing that generated candidates provide complementary value. This is especially important in scenarios where the retrieval database lacks relevant examples, and generation can introduce novel, informative candidates that enhance the overall candidate pool.
@@ 4.4 Hyperparameter Analysis
Figure 3: Hyperparameter analysis at the region level on IM2GPS3K. Trends observed at the region level are representative across different geographic levels. Results for all hyperparameters across all levels can be found in Appendix H.
To better understand the impact of key hyperparameters, we conduct a systematic ablation study by varying each at a time while keeping others fixed. The hyperparameters considered include: (1) Number of retrieval candidates in training $\rvert\mathcal{C}_{{rc}}\rvert$. (2) Number of retrieval candidates during inference $\rvert\mathcal{C}_{{r}}\rvert$. (3) Number of generated candidates during inference $\rvert\mathcal{C}_{{g}}\rvert$. (4) Weighting coefficient $\lambda$. (5) Number of top elements involved in the first-order loss ($K^{(1)})$. Unless otherwise specified, we use the following default values: $\rvert\mathcal{C}_{{rc}}\rvert=5$, $\rvert\mathcal{C}_{{r}}\rvert=10$, $\rvert\mathcal{C}_{{g}}\rvert=5$, $\lambda_{1}=0.7$, and $K^{(1)}=1$. Empirically, we find that the trends of each hyperparameter remain largely consistent across all levels, indicating the stability and robustness of our model under varying geolocalization granularities. For clarity and brevity, we present results at the region level as a representative example in Figure 3, while full results for all levels are provided in Appendix H. In addition, an ablation study on the retrieval pool size is provided in Appendix I for reference.
@@ 4.7 Impact of Backbone Scale
To investigate the impact of backbone model scale on performance, we conduct experiments using llava-onevision li2024llava (0.5B) and Qwen2-VL models wang2024qwen2 with 2B and 7B parameters. As shown in Figure 6, and results across all geographic levels in Appendix J, GeoRanker’s performance consistently improves as the backbone model size increases on both IM2GPS3K and YFCC4K. These results indicate that GeoRanker benefits from more powerful LVLM backbones and follows the scaling law, suggesting that its upper-bound performance can be further improved with larger models.
@@ Acknowledgements
The authors would like to thank Leitian Tao for the valuable feedback on the work. Pengyue Jia and Xiangyu Zhao are supported by the Institute of Digital Medicine of City University of Hong Kong (No.9229503), Research Impact Fund (No.R1015-23), Collaborative Research Fund (No.C1043-24GF), and General Research Fund (No.11218325).
@@ Appendix D Baseline Method Details
  • $\text{[L]kNN},\sigma=4$ [1]. kNN first retrieves the top-$k$ nearest neighbor images and aggregates their coordinates to form the final prediction. As the $k$ decreases, the aggregation process becomes more focused. When $k$ equals 1, the method turns to the NN.
@@
  • CPlaNet [15]. CPlaNet follows PlaNet and proposes combinatorial partitioning, which generates fine-grained output classes by intersecting larger partitions.
@@
  • Img2Loc [10]. Img2Loc advances geolocalization by integrating an RAG pipeline. It first retrieves visually similar candidates, then formulates a prompt incorporating these candidates’ coordinates, guiding a vision-language model to generate a final prediction.
@@
  • G3 [14]. G3 proposes a three-stage framework comprising Geo-alignment, Geo-diversification, and Geo-verification. Geo-alignment aligns GPS coordinates, textual descriptions, and visual data into a unified multi-modal representation to strengthen retrieval capabilities. Subsequently, Geo-diversification and Geo-verification are integrated within an RAG framework to robustly generate and select candidate geolocations.
@@ Appendix E Query Images with Different Error Thresholds
Appendix F Experimental Analysis with Ranking Metrics
Table 5: Ranking performance comparison on IM2GPS3K.
Methods Recall@1 Recall@5 Recall@10 NDCG@5 NDCG@10 NDCG@20
G3 Retrieval 0.0894 0.3217 0.5672 0.6238 0.6735 0.7989
GeoRanker 0.1982 0.5169 0.7387 0.8026 0.8419 0.8872
To comprehensively assess the ranking capability of GeoRanker as a ranking model, we additionally adopt standard learning-to-rank metrics, specifically Recall@K and NDCG@K. We take the top-20 candidates retrieved by G3 as the candidate pool. For Recall@K, we designate the most accurate candidate (i.e., the one closest to the ground truth) as the ground truth candidate and compute whether the ground truth candidate appears within the top-k predictions. For NDCG@K, we assign relevance scores based on distance to the ground truth: candidates within 1 km are assigned a label of 1.0; those within 1–25 km receive 0.8; 25–200 km receive 0.6; 200–750 km receive 0.4; 750–2500 km receive 0.2; and those beyond 2500 km are labeled 0. The results are shown in Table 5: GeoRanker significantly improves both Recall and NDCG over G3. This demonstrates that our ranking module effectively re-orders the candidates to improve fine-grained geolocalization accuracy.
Appendix G Complete experimental results on ablation study